OpenStudy (anonymous):

Find the distance between the points (-4, 6) and (8, -12).Round your solution to 2 decimal points.

OpenStudy (lgbasallote):

use distance formula \[\large \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] \[\large \implies \sqrt{(-4 - 8)^2 + \left[6-(-12)\right]^2}\] does that help?

OpenStudy (anonymous):

yes, so i solve the equation by doing the equation and then the root?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks